Trace the conicoid represented by x
2 +2z
2 = y. Also describe its sections by the
planes x = c,∀c ∈ R.
1
Expert's answer
2020-03-21T17:17:30-0400

"x^2+2z^2=y"  is a paraboloid with vertex at (0, 0, 0) opening away from the origin centered on the line "x=0, z=0".

The conicoid cross section with planes perpendicular to its centered axis are ellipses with the formula "\frac {x^2}{2}+z^2=\frac{y}{2}"(fig.1 ) The conicoid is the Elliptical paraboloid (fig.2).

When "y=0", "x^2+2z^2=0" , which means , so that this section is a single point, (0, 0, 0).

When "x=0""y=2z^2" which is a parabola in the "YZ" plane with vertex at (0, 0, 0).

When "z=0, y=x^2" , and this section is a parabola in the "YX" plane with vertex at (0, 0, 0).

The sections of a paraboloid by the planes "x = c,\u2200c \u2208 R" .

are parabolas "y=2z^2+c^2" which are shifted as a whole in the direction of large values of "y" if "|c|" is increased. For "x=c" and "x=-c" parabolas, in the corresponding planes are the same.

fig.1


fig.2